OpenStudy (anonymous):

2a^2-4a+2/3a^2-3 Please show all work, and state any excuded values.

OpenStudy (luigi0210):

factor

OpenStudy (anonymous):

idk how

OpenStudy (luigi0210):

Well we are screwed now aren't we? :P

OpenStudy (anonymous):

Ok this is what i got: 2a^2-4a+2/3a^2-3=2(a-1)/3(a+1)=\[a \neq-1\]

OpenStudy (anonymous):

But my teacher said I need to show my work so now what? @Hero

OpenStudy (anonymous):

@mathstudent55

OpenStudy (anonymous):

@a

OpenStudy (anonymous):

@anikhalder @abbiebosworth @aditkarekatte @ariannafar1 @alli14344 @Ashleyisakitty @bobsr. @countrygirl33 @dumbcow @dan815 @deadman340

OpenStudy (luigi0210):

@Jhannybean she'll take care of you

OpenStudy (jhannybean):

omg lol he tagged.....a lot of people.

OpenStudy (anonymous):

lol

OpenStudy (anikhalder):

the second a^2 in 2/3a^2 is in the denominator?

OpenStudy (jhannybean):

\[\large \frac{2a^2-4a+2}{3a^2-3}\] first step, let's start with the numerator. divide the numerator by 2. what do you get?

OpenStudy (luigi0210):

wait for it...

OpenStudy (jhannybean):

second step, divide denominator by 3, what do you get?

OpenStudy (anonymous):

Um idk 3a^2?

OpenStudy (luigi0210):

\[\frac{ 2(a^2-2a+1) }{ 3(a^2-1) }\]

OpenStudy (jhannybean):

lol not factor OUT, divide!!!

OpenStudy (luigi0210):

I'm not sure if you can divide..

OpenStudy (luigi0210):

but then again you're the expert

OpenStudy (jhannybean):

numerator : (2a^2)/2 - (4a)/2 + ( 2 )/2 denominator (3a^2)/3 - (3)/3

OpenStudy (jhannybean):

can you simplify these? lol. That's what i meant.

OpenStudy (anonymous):

ok 2(a-1)/3(a+1)?

OpenStudy (luigi0210):

see you can factor, good job

OpenStudy (luigi0210):

forgot to factor that 3 Jhann

OpenStudy (anonymous):

yea but thts wht i did and then i put a doesnt equal -1 but my teacher said i need to show my work...WTF

OpenStudy (jhannybean):

OMGGGG..

OpenStudy (anonymous):